# Human-Like Robot Impedance Regulation Skill Learning from Human-Human Demonstrations

### v1

@@ III Methodology
In this section, the overall framework (as shown in Fig. 3) will be presented. The interaction model is introduced to represent the dynamics between the agent and the environment. The endpoint stiffness matrices along the trajectories are estimated and the reference trajectories of the demonstrator are then calculated with data recorded from human demonstrations. A novel TP-GMM is proposed for learning motion and impedance representation from the demonstrations by corresponding the reference trajectory and the stiffness profiles. The LQT method for trajectory planning and impedance variation is then raised afterwards for robot execution. Meanwhile, an impedance regulation module to construct the compliance relationship between two demonstrators is proposed. Finally, a whole-body impedance controller is applied to perform the reference trajectory and impedance variables for the CURI.
@@ III-A Interaction Model
The interaction model between the cobot and the environment encapsulates the dynamics of the physical behaviors during the collaboration, which is supposed to imitate human collaboration skills by utilizing both trajectory and force information. Hence, we propose to learn the reference trajectory and the stiffness matrix of the interaction model between the human and the environment during the human-human demonstrations and then transfer to the cobot for human-robot collaboration.
Considering the movement of the cobot Tool Center Point (TCP) as a single unit mass moving in Cartesian space under the effect of a control input and interaction force. To achieve the desired dynamics of the tasks, the problem can be formulated as designing the appropriate control input. Specifically, the Cartesian-space impedance model to dynamically connect the cobot positions, velocities and accelerations with the interactive forces is:
$\bm{\Lambda}^{d}\bm{\ddot{\widetilde{x}}}+\bm{D}^{d}\bm{\dot{\widetilde{x}}}+\bm{K}^{d}\bm{\widetilde{x}}=\bm{F}^{ext},$ (7)
where $\bm{\widetilde{x}}=\bm{x}^{d}-\bm{x}\in\mathbb{R}^{6}$ in which $\bm{x}^{d}$ is the reference trajectory of the virtual spring-damper system and $\bm{x}$ the TCP pose. $\bm{\Lambda}^{d},\bm{D}^{d},\bm{K}^{d}\in\mathbb{R}^{6\times 6}$ are the symmetric and positive definite matrices of the desired inertia, damping, and stiffness, respectively. $\bm{F}^{ext}$ is the external force applied to the end effector.
As mentioned above, we propose to learn the reference trajectory for cobot to execute. However, we can only observe the actual trajectory $\bm{x}$ of the demonstrations instead of the variable $\bm{x}^{d}$. Assuming the velocity and acceleration of the reference trajectory remain at 0 [5], the dynamic model can be rewritten based on (7):
$\bm{\ddot{x}}=\bm{K}^{d}(\bm{x}^{d}-\bm{x})-\bm{D}^{d}\bm{\dot{}}{{x}}+\bm{F}^{ext}.$ (8)
The reference trajectory $\bm{x}^{d}$ can be calculated accordingly if endpoint pose $\bm{x}$, stiffness (damping) matrix $\bm{K}^{d}$ ($\bm{D}^{d}$), and external forces $\bm{F}^{ext}$ are known.
@@ III-B Motion Representation and Planning
The reference trajectory of each demonstration is firstly calculated by referring to (8). The newly generated trajectories can be observed from different frames and GMM is adopted to cluster points on the demonstration into multiple Gaussian distributions in each frame $i\in[1,2]$. Compared with previous works that applied TP-GMM [12], our method binds motion and impedance profiles and then calculates the GMM in multiple dimensions (7 for pose and 3 for impedance). After that, a motion with a corresponding impedance profile considered the task parameter can be reproduced via the product of experts (PoE). The task parameter refers to the pose of the target object or the state of the environment, which depends on the specific task. Note that the task parameter is defined as the transformation matrices of the start point and endpoint of multiple demonstrations $\bm{\zeta}$ from the world origin, namely $\{\bm{b}^{(i)},\bm{A}^{(i)}\}_{i=1,2}$. We first use these matrices to transform and align demonstrations with respect to each observer frame $\bm{X}_{t}^{(i)}={\bm{A}^{(i)}}^{-1}(\bm{\zeta}_{t}-\bm{b}^{(i)})$. Then we use a TP-GMM with $J$ Gaussian distributions $\{\bm{\mu}_{j}^{(i)},\bm{\Sigma}_{j}^{(i)}\}_{i=1,2},j\in[1,J]$ to represent $\bm{X}_{t}^{(i)}$ in each frame. Thus, the reproduction is expected to lie within the distributions $\mathcal{N}\left(\hat{\bm{\xi}}^{(i)}_{j},\hat{\bm{\Sigma}}^{(i)}_{j}\right)_{i=1,2}$, where $\hat{\bm{\xi}}^{(i)}_{j}=\bm{A}^{(i)}\bm{\mu}^{(i)}_{j}+\bm{b}^{(i)}$, $\hat{\bm{\Sigma}}^{(i)}_{j}=\bm{A}^{(i)}\bm{\Sigma}^{(i)}_{j}{\bm{A}^{(i)}}^{\top}$.
@@
$\hat{\bm{\xi}}$ $=\arg\min_{\bm{\xi}}\sum_{j=1}^{2}\left(\hat{\bm{\xi}}_{k,t}^{(j)}-\hat{\bm{\zeta}}\right)^{\top}{\hat{\bm{\Sigma}}_{k,t}^{(j)^{-1}}}\left(\hat{\bm{\xi}}_{k,t}^{(j)}-\hat{\bm{\zeta}}\right),$ (11)
@@ III-C Impedance Regulation Skill Learning
where $\bm{V}^{c}$, $\bm{V}^{p}$, $\bm{H}_{s}^{c}$, and $\bm{H}_{s}^{c}$ represent the orthonormal matrices and the diagonal matrices containing the configuration dependence eigenvalues of the collaborator and the partner, respectively. The arm configurations of two individuals remain corresponding during human demonstrations, which makes the variation of $\bm{V}$ and $\bm{H}_{s}$ basically unchanged along the movement from one demonstration to another. Hence, the variation of the regulatory factor during tasks will be learned between the two synergistic contributions of the muscle co-contractions $\alpha(A^{c})$ and $\alpha(A^{p})$.
Machine learning has a high potential to build the implicit model between two individuals in collaboration, especially for those tasks with periodicity [34]. We propose to use a Long-Short-Term-Memory (LSTM) neural network for regulation policy learning since the data of muscular synergistic contributions are strictly based on time series during the demonstrations [35]. Specifically, the data will constitute the form with the LSTM function as $h:X\rightarrow Y$. A loss function will be defined to evaluate the performance of the model by comparing the predicted outputs with the actual results.
The synergistic contribution values of the partner and collaborator are chosen as the inputs and outputs of the network model, respectively. LSTM is sensitive to the scale of the input data, especially when using the sigmoid activation function. Hence, we normalize the raw data to a range of $[0,1]$ as inputs and reverse the predictive outputs from the model for better training. Note that 67% of the demonstration data will be used for training while the remaining 33% will be used for testing. We connect multiple sets of demonstration data of one specific task end to end to form the dataset for model construction. One example of the model training and testing results was given in the upper part of Fig. 5, which shows the comparison between the predicted muscle synergistic contributions and the ground truth values of the collaborator during object transportation demonstration. We applied Root Mean Squared Error (RMSE) to evaluate the accuracy of predictions with the LSTM network. The prediction performance of the testing set (RMSE: 8.98) is basically consistent with that of the training set (RMSE: 8.22).
@@ III-D Whole-Body Impedance Controller
Once the robot dynamic (14) and impedance model (7) are given, the Cartesian impedance controller proposed in [36] can be extended to CURI, while its stability can also be guaranteed. The whole-body impedance controller generating high-level torque command for each chain $\tau_{i}$ ($i=l,r$) can be expressed as:
@@
where $\bm{\tau}_{i}^{0}$ represents the corresponding torque projected in the null space of the main task, $\bm{N}(\bm{q}_{i})=\bm{I}-\bm{J}_{i}^{T}\bm{J}_{i}^{+T}\in\mathbb{R}^{10\times 10}$ is the projection matrix in order to prevent interference with the Cartesian impedance behavior of the main task according to [37]. $\bm{J}_{i}^{+}$ expresses the generalized inverse of $\bm{J}_{i}$.
@@
The above formulation allows us to control one chain of CURI at a time. However, it may cause conflicts when we are supposed to control both chains at the same time since the torso are shared parts. To address this problem, we propose a leader-follower method regarding the two chains. Assuming the left arm as leader-arm, the whole-body impedance controller for its corresponding chain can be designed with $\bm{\tau}_{l}^{0}$ in (17), which is generated by some predefined sub-tasks, i.e., optimizing the joint configuration for the robotic arm [23]. This paper defines the sub-task as the torso movement strategy based on human demonstrations. Specifically, the cobot torso is supposed to move in the same direction as the demonstrator’s trunk at different stages of the task [38]. Then the calculated $\bm{\tau}_{t}$ and $\bm{\tau}_{al}$ of the leader-arm can be treated as a sub-task of the controller corresponding to the right chain of CURI, whose corresponding arm is named as follower-arm. The $\bm{\tau}_{r}^{0}$ of follower-arm can be chosen as:
[…]
This impedance model allows us to change the cobot physical behavior by varying both the damping and stiffness while tracking the reference trajectory according to the task requirements. In order to avoid instability of the impedance control method due to arbitrary variations in impedance parameters, some passivity conditions of the varying impedance parameters should be satisfied to guarantee system stability [39]. Under the circumstances that the conditions are met, the calculated impedance variables $\bm{K}^{d}_{r}$ and $\bm{D}^{d}_{r}$ from cobot reference trajectory and stiffness generation module can be fed into the controller to perform expected physical behaviors.
@@ IV Experiments
Two typical tasks were introduced to verify the performance of the proposed learning framework under leader-follower mode and mutual adaptation mode, respectively. The comparative experiments of the proposed variable impedance control method and a constant impedance control method have illustrated the advantages of our method.
@@ IV-A Endpoint Stiffness Parameter Calibration
We designed perturbation experiments to identify the parameters of each demonstrator, which will be further applied to extract the endpoint stiffness during human-human demonstrations. The experimental setup is illustrated in Fig. 6. We used one arm of the CURI for parameter identification, which includes a force/torque (F/T) sensor (SRI, 100 $Hz$) connecting the handle and the end effector. An optical system (Optitrack, 120 $Hz$) was applied to track the poses of the shoulder, elbow, and wrist by attaching markers to these human parts. Besides, a wireless EMG device (Delsys Trigno, 2000 $Hz$) and a monitor were used to visualize the real-time muscle co-contraction.
We chose BB and TB on the upper arm as antagonistic muscles to calculate the co-contraction level as mentioned in Section II-B. Specifically, the raw EMG signals were collected and processed using a lowpass filter with a frequency of 50 $Hz$ to eliminate noise. Then the activation level of each muscle was normalized by MVC. Note that the values of each muscle MVC were acquired before the experiment. Finally, the co-contraction level was calculated based on (6).
Figure 6: Experimental setup for endpoint stiffness parameters calibration. Six different human arm configurations $a-f$ were selected for calibration through physical human-robot interaction. (We declare that the individual in this figure is author and he gave permission for the use of his image.)
During perturbation experiments, a trajectory with perturbations in six directions ($x$, $y$, $z$, $xy$, $xz$, $yz$) was given for each test. The order in which these perturbations appear was randomized to avoid subject adaptation. The peak displacement of each perturbation in one trial was set as 0.02 $m$ while the duration was 0.5 $s$. We have tested six arm configurations with three different muscle co-contraction levels of each configuration, which are minimum co-contraction, 20% MVC, and 40% MVC. Note that we applied Biodex system to get the MVC values of the target muscles. Hence, the total number of trials for each subject was 18 during stiffness parameters calibration. The parameters ${a_{1},a_{2},b_{1},b_{2}}$ of two subjects in this paper were given as follows in Table I.
@@
Subject 1 0.272 1.314 3847.141 151.684
Subject 2 0.107 2.200 2678.765 149.597
Figure 7: Data collection setup: Two individuals with endpoint stiffness calibration were attached to the Optitrack markers on their arm joints while the co-contraction muscle signals were recorded by EMG sensors. F/T sensors were attached to the object. The snapshots of human-human demonstration. I: collaborative transportation, II.A: Tai Chi unimanual flat circle pushing, II.B: Tai Chi bimanual vertical circle pushing. (We declare that the individuals in this figure are both authors and they gave permission for the use of their image.)
Figure 8: Learning results for motion/impedance representation and generation by TP-LQT method are illustrated. GMM is firstly adopted in coordinates 1 and 2 to cluster points into multiple Gaussian distributions and represent the demonstrations with PoE (a-c). By considering the task parameter with a new situation, a trajectory is generated via PoE (d). Finally, the corresponding stiffness ellipses are drawn along the generated trajectory in $xy$ and $xz$ plane with the LQT method (e-f). I: collaborative transportation, II.A: Tai Chi unimanual flat circle pushing.
IV-B Collaborative Transportation Task

### v2

@@ III Methodology
In this section, the overall framework (as shown in Fig. 4) will be presented. We first introduce an interaction model that represents the contact dynamics between the robot end-effector and the environment, where ‘environment’ denotes the external physical counterparts in contact during the specific task (e.g., manipulated objects and/or the human collaborator). The endpoint stiffness matrices along the endpoint trajectories are estimated and the reference trajectories are then calculated with data recorded from human demonstrators. A TP-GMM is proposed for learning motion and impedance representation from the demonstrations by corresponding the reference trajectory and the stiffness profiles. The LQT method for trajectory planning and impedance variation is then raised afterwards for robot execution. Meanwhile, an impedance regulation module to learn the compliance relationship between two demonstrators is proposed. Finally, a whole-body impedance controller is applied to perform the reference trajectory and impedance variables for the CURI.
@@ III-A Interaction Model
The interaction model between the cobot and the environment encapsulates the dynamics of the physical behaviors during the collaboration, which is supposed to imitate human collaboration skills by utilizing both the Tool Center Point (TCP) trajectory $\bm{x}^{c}$ and force information $\bm{F}^{c}_{ext}$. Hence, we propose to learn the reference trajectory $\bm{x}^{c}_{d}$ and the stiffness matrix of the interaction model between the human and the environment during the human-human demonstrations and then transfer to the cobot for HRC.
Considering the movement of the robot/human TCP ($\bm{x}^{r}$$\bm{x}^{c}$, represented as $\bm{x}$ in the following formulation) as a point-mass moving in Cartesian space under the effect of a control input and interaction force. To achieve the desired dynamics of the tasks, the problem can be formulated as designing the appropriate control input. Specifically, the Cartesian-space impedance model to dynamically connect the positions, velocities and accelerations with the interactive forces is:
$\bm{\Lambda}_{d}\bm{\ddot{\widetilde{x}}}+\bm{D}_{d}\bm{\dot{\widetilde{x}}}+\bm{K}_{d}\bm{\widetilde{x}}=\bm{F}_{ext},$ (7)
where $\bm{\widetilde{x}}=\bm{x}_{d}-\bm{x}\in\mathbb{R}^{6}$. $\bm{\Lambda}_{d},\bm{D}_{d},\bm{K}_{d}\in\mathbb{R}^{6\times 6}$ are the symmetric and positive definite matrices of the desired inertia, damping, and stiffness, respectively. $\bm{F}_{ext}$ is the external force applied to the TCP.
As mentioned above, we propose to learn the reference trajectory for the cobot to execute. However, we can only observe the human TCP trajectory $\bm{x}^{c}$ of the demonstrations instead of the variable $\bm{x}^{c}_{d}$. Assuming the velocity and acceleration of the reference trajectory remain at 0 [rozo2016learning], the dynamic model can be rewritten based on (7):
$\bm{\ddot{x}}=\bm{K}_{d}(\bm{x}_{d}-\bm{x})-\bm{D}_{d}\bm{\dot{}}{{x}}+\bm{F}_{ext}.$ (8)
The reference trajectory $\bm{x}_{d}$ can be calculated accordingly if TCP pose $\bm{x}$, stiffness (damping) matrix $\bm{K}_{d}$ ($\bm{D}_{d}$), and external forces $\bm{F}_{ext}$ are known.
@@ III-B Motion Representation and Planning
The reference trajectory of each demonstration is firstly calculated by referring to (8). The newly generated trajectories can be observed from different frames and GMM is adopted to cluster points on the demonstration into multiple Gaussian distributions in each frame $i\in[1,2]$. Compared with previous works that applied TP-GMM [calinon2016tutorial], our method binds motion and impedance profiles and then fits a joint GMM over a concatenated feature space (7D pose ++ 3D stiffness). After that, a motion with a corresponding impedance profile considered the task parameter can be reproduced via the product of experts (PoE). The task parameter denotes the task-specific frames that condition the TP-GMM. For box transportation, the start/goal poses (and induced height/distance) capture new start–goal configurations. For Tai Chi pushing hands, partner-aligned frames derived from each collaborator’s kinematics capture new collaborator reach and workspace. Note that the task parameter is defined as the transformation matrices of the start point and endpoint of multiple demonstrations $\bm{\zeta}$ from the world origin, namely $\{\bm{b}^{(i)},\bm{A}^{(i)}\}_{i=1,2}$. We first use these matrices to transform and align demonstrations with respect to each observer frame $\bm{X}_{t}^{(i)}={\bm{A}^{(i)}}^{-1}(\bm{\zeta}_{t}-\bm{b}^{(i)})$. Then we use a TP-GMM with $J$ Gaussian distributions $\{\bm{\mu}_{j}^{(i)},\bm{\Sigma}_{j}^{(i)}\}_{i=1,2},j\in[1,J]$ to represent $\bm{X}_{t}^{(i)}$ in each frame. Thus, the reproduction is expected to lie within the distributions $\mathcal{N}\left(\hat{\bm{\xi}}^{(i)}_{j},\hat{\bm{\Sigma}}^{(i)}_{j}\right)_{i=1,2}$, where $\hat{\bm{\xi}}^{(i)}_{j}=\bm{A}^{(i)}\bm{\mu}^{(i)}_{j}+\bm{b}^{(i)}$, $\hat{\bm{\Sigma}}^{(i)}_{j}=\bm{A}^{(i)}\bm{\Sigma}^{(i)}_{j}{\bm{A}^{(i)}}^{\top}$.
@@
[…]
@@ III-C Impedance Regulation Skill Learning
where $\bm{V}^{c}$, $\bm{V}^{p}$, $\bm{H}_{s}^{c}$, and $\bm{H}_{s}^{p}$ represent the orthonormal matrices and the diagonal matrices containing the configuration dependence eigenvalues of the collaborator and the partner, respectively. The arm configurations of two individuals remain corresponding during human demonstrations, which makes the variation of $\bm{V}$ and $\bm{H}_{s}$ basically unchanged along the movement from one demonstration to another. Hence, the variation of the regulatory factor during tasks will be learned between the two synergistic contributions of the muscle co-contractions $\alpha(A^{c})$ and $\alpha(A^{p})$.
Machine learning has a high potential to build the implicit model between two individuals in collaboration, especially for those tasks with periodicity [mukherjee2022survey]. We propose to use a Long-Short-Term Memory (LSTM) neural network for regulation policy learning since the data of muscular synergistic contributions are strictly based on time series during the demonstrations [semeraro2023human]. These kinds of time series prediction models have been applied to model the interaction behaviors, such as EMG-to-force [xie2024deep] and human intention recognition [9568762, schirmer2023]. To achieve this, the data will constitute the form with the LSTM function as $h:X\rightarrow Y$. A loss function will be defined to evaluate the performance of the model by comparing the predicted outputs with the actual results.
In this paper, the LSTM network model is trained using input data consisting of the partner’s muscle co-contraction signals and the corresponding demonstrated trajectories, while the output is designed to predict the collaborator’s muscle co-contraction. Hence, the form with the LSTM function as $h:(\alpha(A^{p}),\bm{\zeta})\rightarrow\alpha(A^{c})$. This design enables the model to capture the temporal relationships between partner behaviors and task dynamics. By learning this mapping, the trained model can generalize to different trajectories, which may arise due to variations in individual personalities or movement styles, thereby enhancing its adaptability to diverse interaction scenarios. LSTM is sensitive to the scale of the input data, especially when using the sigmoid activation function. Therefore, we normalize the raw data to a range of $[0,1]$ as inputs and de-normalize the predictive outputs from the model for better training. We connect multiple sets of demonstration data end-to-end to form the dataset for model construction.
The effectiveness of this design is validated in Section IV, where the proposed LSTM-based model demonstrates its capability to adapt to various interaction scenarios. Specifically, the experiments show that the model successfully generalizes to diverse trajectories, accommodating differences in individual movement styles.
@@ III-D Whole-Body Impedance Controller
Once the robot dynamic (14) and impedance model (7) are given, the Cartesian impedance controller proposed in [ott2008cartesian] can be extended to CURI, while its stability can also be guaranteed. The whole-body impedance controller generating high-level torque command for each chain $\tau_{i}$ ($i=l,r$) can be expressed as:
@@
Figure 5: Calibration setup: endpoint stiffness parameters calibration through perturbation methods. Data collection setup: two subjects were attached to the OptiTrack markers on their arm joints while the co-contraction muscle signals were recorded by EMG sensors. F/T sensors were attached to the object/subject. Perturbation experiments: six different human arm configurations afa-f were selected for calibration through physical human-robot interaction. The snapshots of human-human demonstrations. I: collaborative transportation, II.A: Tai Chi unimanual flat circle pushing, II.B: Tai Chi bimanual vertical circle pushing, and III: collaborative sawing.
@@
where $\bm{\tau}_{i}^{0}$ represents the corresponding torque projected in the null space of the main task, $\bm{N}(\bm{q}_{i})=\bm{I}-\bm{J}_{i}^{T}\bm{J}_{i}^{+T}\in\mathbb{R}^{10\times 10}$ is the projection matrix in order to prevent interference with the Cartesian impedance behavior of the main task. $\bm{J}_{i}^{+}$ expresses the generalized inverse of $\bm{J}_{i}$.
@@
The above formulation allows us to control one chain of CURI at a time. However, it may cause conflicts when we are supposed to control both chains at the same time since the torso are shared parts. To address this problem, we propose a leader-follower method regarding the two chains. Assuming the left arm as the leader arm, the whole-body impedance controller for its corresponding chain can be designed with $\bm{\tau}_{l}^{0}$ in (17), which is generated by some predefined sub-tasks, i.e., optimizing the joint configuration for the robotic arm [dietrich2016whole]. This paper defines the sub-task as the torso movement strategy based on human demonstrations. Specifically, the cobot torso is supposed to move in the same direction as the demonstrator’s trunk at different stages of the task [neo2007whole]. Then the calculated $\bm{\tau}_{t}$ and $\bm{\tau}_{al}$ of the leader arm can be treated as a sub-task of the controller corresponding to the right chain of CURI, whose corresponding arm is referred to as the follower arm. The $\bm{\tau}_{r}^{0}$ of the follower arm can be chosen as:
$\bm{\tau}_{r}^{0}=\left(\begin{matrix}\bm{\tau}_{t}^{T},\ \bm{0}_{1\times 7}\end{matrix}\right).$ (18)
This impedance model enables the cobot to perform human-like behavior adjustment by varying damping and stiffness while tracking the reference trajectory based on task requirements. To ensure stability, passivity conditions for time-varying impedance parameters must be satisfied [kronander2016stability]. When these conditions are met, the calculated impedance variables $\bm{K}_{d}^{r}$ and $\bm{D}_{d}^{r}$ can be fed into the controller to achieve the desired behaviors.
@@ IV Experiments
Three scenarios including four typical tasks were introduced to verify the performance of our framework under the leader-follower mode and the mutual adaptation mode, which are shown in Fig. 5. Five human subjects were involved with four males and one female (age: $28\pm 2$ years; height: $1.74\pm 0.14$ $m$). Comparative experiments of our proposed HI-ImpRSL, a fixed impedance control method (FIC) [9362246], an EMG-based variable impedance control method (EMG-based VIC) [peternel2016adaptation], a neural network-based variable impedance skill learning framework (NN-based VIC) with interaction force-based stiffness estimation [zhang2023neural], and a TP-LQT based variable impedance control method (HI-ImpRSL w/o LSTM) [rozo2016learning], illustrate the advantages of our HI-ImpRSL method.
@@ IV-A Endpoint Stiffness Parameter Calibration
Perturbation experiments are designed to identify the parameters of each demonstrator, which are applied to extract the endpoint stiffness during human-human demonstrations. The calibration setup is illustrated in Fig. 5. We used CURI arm for parameter identification and mounted a force/torque (F/T) sensor (SRI, 100 $Hz$) connecting the handle and the end effector. An optical system (OptiTrack, 120 $Hz$) was applied to track the poses of the shoulder, elbow, and wrist by attaching markers to these human joints. Besides, a wireless EMG device (Delsys Trigno, 2000 $Hz$) and a monitor were used to visualize the real-time muscle co-contraction.
We selected biceps brachii (BB) and triceps brachii (TB) as antagonistic pairs to compute co-contraction (Sec. II-B). EMG was acquired with a Delsys Trigno system at 2000 $Hz$, followed by a fourth-order 100 $Hz$ high-pass Butterworth filter to clean the raw data. Signals were full-wave rectified and processed for the envelope at 250 $Hz$, each muscle’s activation was normalized to its MVC obtained with a Biodex system, and activations were scaled to $[0,1]$. The resulting activations were used to compute co-contraction in (6).
During perturbation experiments, a trajectory with perturbations in six directions ($x$, $y$, $z$, xyxy, xzxz, yzyz) was given for each test. The order in which these perturbations appear was randomized to avoid subject adaptation. The peak displacement of each perturbation in one trial was set as 0.02 $m$ while the duration was 0.5 $s$. We have tested six arm configurations with three different muscle co-contraction levels of each configuration, which are minimum co-contraction, 20% MVC, and 40% MVC. Hence, the total number of trials for each subject was 18 during stiffness parameters calibration. The parameters $a_{1},a_{2},b_{1},b_{2}$} of five subjects in this paper were given as follows in Table I.
@@
S1 0.272 1.314 3847.141 151.684
S2 0.107 2.200 2678.765 149.597
S3 0.399 2.926 1819.695 128.581
S4 0.341 4.073 2699.123 112.562
S5 0.167 4.528 1260.290 94.951
Figure 6: Learning results for motion/impedance representation and generation by TP-LQT method are illustrated. GMM is firstly adopted in coordinates 1 and 2 to cluster points into multiple Gaussian distributions and represent the demonstrations with PoE (a-c). By considering the task parameter with a new situation, a trajectory is generated via PoE (d). Finally, the corresponding stiffness ellipses are drawn along the generated trajectory in planes with the LQT method (e-f). I: collaborative transportation, II.A: Tai Chi unimanual flat circle pushing.
IV-B Collaborative Transportation